OpenStudy (anonymous):

A geometric sequence is defined recursively by a(n) = 4a(n-1). The 10th term of the sequence is a(10) = 3.5. What is a(13)?

OpenStudy (amistre64):

4 is the ratio right?

OpenStudy (amistre64):

A{n} = 4A{n-1} = 4(4A{n-2}) = 4(4(4A{n-3})) = 4^3 A{n-3} = 4^r A{n-r} ; when n-r = 0 when r=n = 4^n A{0} A{10} = 3.5 = 4^10 A{0} 3.5/(4^10) = A{0} A{n} = 4^n (3.5)/4^10 ... maybe :)

OpenStudy (amistre64):

A{10} = 4^10 (3.5)/4^10 = 3.5 ; lol A{13} = 4^13 (3.5)/4^10 = 224